Now every column which has no circled entry is completely zero; so when k = 6 and k = 7 the algorithm outputs two more vectors, namely

(1) v^{[2]} = (0, 5, 5, 0, 9, 5, 1, 0) v^{[3]} = (0, 9, 11, 9, 10, 12, 0, 1)

From the form of the matrix A after k = 5, it is evident that these vectors satisfy the equation vA = (0, …, 0).